Opened 21 months ago

Closed 4 months ago

#605 closed flight processing (fixed)

EUFAR15/68, flight day 270/2015, Ain

Reported by: asm Owned by:
Priority: immediate Milestone: 2015 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by dac)

Data location: ~arsf/arsf_data/2015/flight_data/france/EUFAR15_68-2015_270_Ain

Project acronym: HYMOSENS2

Data arrived from ARSF via SATA disk 8 on 28/09/2015.

Priority: Unknown

PI: Luca Demarchi

EUFAR code: E068

Sensors:

  • Fenix (not required)
  • Leica LIDAR (requested)
  • Leica FW (not requested)
  • RCD (requested)
  • Owl (not requested)

Change History (24)

comment:1 Changed 21 months ago by asm

  • Milestone set to The Glorious Future
Last edited 21 months ago by asm (previous) (diff)

comment:2 Changed 21 months ago by asm

  • Milestone changed from The Glorious Future to 2015 data processing completion

comment:3 Changed 19 months ago by dac

From Gary 05/01/2016 "all LiDAR and RCD to be processed. Only Fenix with 2 hours of solar noon to be processed. Speak to Gary for clarification if needed."

comment:4 Changed 19 months ago by dac

  • Description modified (diff)

Fenix data don't need processing for this flight.

comment:5 Changed 18 months ago by lah

Basestation

Had to convert basestation file using tecq even though only 1 survey in file.

Lat: 45 43 18.26917
Lon: 4 56 37.29527
Elipsoidal Height: 251.029 (metres)

comment:6 Changed 18 months ago by lah

Navigation processing complete

comment:7 Changed 18 months ago by lah

RCD Processing

  • Converted to Tiffs
  • moved lidar log files into correct directory

now tagging

comment:8 Changed 18 months ago by lah

RCD Delivery created

  • set event file group read permissions (script defaults to airborne only)

ready for delivery check

comment:9 Changed 18 months ago by dac

RCD Delivery Check

Started delivery check.

comment:10 Changed 18 months ago by dac

RCD

Delivery looks fine. Ready to deliver once payment is received from BAS.

comment:11 Changed 17 months ago by dac

LiDAR

Started processing.

comment:12 Changed 17 months ago by dac

PI emailed to check delivery preference.

comment:13 Changed 17 months ago by dac

LiDAR

Good match using boresight pitch and roll. Classified discrete files - copying to full waveform.

comment:14 Changed 17 months ago by dac

LiDAR

Delivery created - waiting for delivery check,

comment:15 Changed 17 months ago by asm

Lidar Delivery Checks

All checks have been done and everything is OK except demcompare. Script is throwing an error and still not sure which development branch has solved problems encountered. Once checked, should be ready to deliver.

Also prject_tidy complains about name of logsheet (EUFAR15_68 instead of EUFAR_15_68, already changed)

Last edited 17 months ago by asm (previous) (diff)

comment:16 Changed 17 months ago by asm

Lidar DC

Demcompare working now. Output is correct.

Difference statistics:
Min: -42.9587644042969
Max: 36.5832888183594
Sum: -96057.1497435304
Mean: -1.91834221523636
Median: -1.31263
Absolute mean: 5.07110954972321
Std deviation: 6.68433221056402
Total cells: 3291750
Non-null cells: 50073

Lidar is ready to deliver

comment:17 Changed 17 months ago by dac

Delivery

PI Requested delivery by hard drive as well as FTP.

comment:18 Changed 17 months ago by asm

Delivery

Upon request from PI, LiDAR and camera data (no Fenix on this project) has also been dispatched on a hard drive via postal service on 17/03/2016.

comment:19 Changed 16 months ago by dac

Archiving

Started uploading processed data to NEODC.

comment:20 Changed 16 months ago by dac

Archiving

Processed data uploaded to NEODC.

comment:22 Changed 14 months ago by dac

Archiving

Started uploading raw data to NEODC.

comment:23 Changed 4 months ago by asm

Archiving

Data now available from NEODC. Closing ticket.

comment:24 Changed 4 months ago by asm

  • Resolution set to fixed
  • Status changed from new to closed
Note: See TracTickets for help on using tickets.